# Resolvi: A Reference Architecture for Scalable and Interoperable Entity Resolution

### v1

@@ Abstract
Entity resolution (ER) is a critical task in data management which identifies whether multiple records refer to the same real-world entity. Despite its significance across domains such as healthcare, finance, and machine learning, implementing effective ER systems remains challenging due to the abundance of methodologies and tools, leading to a paradox of choice for practitioners. This paper proposes Resolvi, a reference architecture aimed at enhancing extensibility, interoperability, and scalability in ER systems. By analyzing existing ER frameworks and literature, we establish a structured approach to designing ER solutions that address common challenges. Additionally, we explore best practices for system implementation and deployment strategies to facilitate large-scale entity resolution. Through this work, we aim to provide a foundational blueprint that assists researchers and practitioners in developing robust, scalable ER systems while reducing the complexity of architectural decisions.
@@ 1 Introduction
The terminology surrounding ER is diverse. It is commonly referred to as record linkage, deduplication, merge/purge, or entity matching [12, 49, 6]. In some contexts, particularly in natural language processing, it appears under the names entity alignment or named entity recognition and disambiguation (NERD) [54]. This terminological variation reflects the breadth and complexity of the field, creating challenges in unifying methodologies and system architectures.
@@ 3.1 Business Requirements
The first such requirement is that it must determine whether multiple information refer to the same real world object. With the risk of stating the obvious, this entails that the system should discern what a real world object is. Human assistance in this regard is arguably required at this point in our technological development. It logically follows that completely unsupervised entity resolution solutions are more unreliable than supervised ones. Consequently, this reference architecture focuses on an ER process which allows humans in the loop at design time.
@@ 3.2 Architectural Characteristics
Just like there are business requirements which determine whether a system performs entity resolution or not, there are architectural characteristics which are generally required of all systems which perform end-to-end entity resolution.
ER is a highly polymorphic task with a huge number of applications. Among the top priorities when implementing any such system is its extensibility. Extensibility is exhibited to various degrees in practice today. Some systems allow using an arbitrary number of data sources [58, 51]. Others can swap out their runtime completely [28, 44].
ER is valuable standalone or as part of a more complex process. Therefore, the best ER systems will be highly interoperable either by supporting many data contracts or by using widely adopted technologies and data formats. Interoperability is an obvious concern to some of the reference implementations [51, 21, 58, 34] that were inspected.
Basic ER implementations that rely on similarities between entity references have a minimum time complexity of $\mathcal{O}(n^{2})$ [12] where $n$ is the total number of entity references available to the system. The high complexity coupled with the need to process large amounts of data makes performance and scalability indispensable architectural characteristics of any modern ER system. Performance is the ability of a system to use the available hardware resources efficiently [53] (i.e the system does not waste computations). Scalability is the ability of a system to use hardware resources effectively as the demand on the system increases [53] (e.g the system is able to access more resources as it receives more input).
@@ 4.2 Entity Resolution Runtime
Some ER systems store entity references and entity profiles for later use. Storing the state is particularly important for perform progressive ER [58, 51, 21] or to preserve stable identifiers across time for entity profiles [68].
@@ 4.3 System Configuration
In existing systems the end user can customize the runtime itself by choosing a default runtime [34, 58] or even building one from existing components exposed by the system [25, 51]. Another runtime customization is the ability to train a configuration optimised for a specific dataset either via active learning [66] or genetic algorithms [21]. This configuration can then be attached to many ER runtime executions. This approach stands in contrast to building the optimal configuration at runtime through unsupervised learning [34]. In that case a runtime configuration of the data record attributes that are relevant in ER is sufficient.
System configuration is exhibited in existing ER systems through plugin management for some [44] or for all of the ER runtime components [58, 51]. Some systems are designed as frameworks [25, 28], enabling the configuration of the runtime architecture itself. For example, [44] allows using neural networks throughout the entire runtime or only part of it. System configuration also refers to configuring the types of data sources or the execution engines supported by the ER system. Some systems [21, 51, 58] even allow choosing how they represent entity references internally.
@@
Specific ER system architectures should involve administrators in most configuration use cases. This recommendation assumes that administrators are better prepared on security and configuration topics than end users and that in practice they usually have broader access to the system than developers. Developers also play a role in the design of the ER runtime. They may train machine learning models or implement plugins that perform various ER tasks such as matching, clustering or data ingestion. These models and plugins are visible to end users through configuration. Developers may register and unregister the models and plugins they author. They also typically maintain plugin implementations and models. Plugin management may be performed by all three roles: administrators, end users and developers. End users might focus on choosing the plugins they need to configure custom runtimes. Developers would focus on managing their own plugins. Finally, administrators could enforce system-wide rules for using plugins. A system architecture that allows external developer involvement (through plugin management or otherwise), may be desirable. However, it raises many potential concerns, primarily from the sphere of data privacy and security. Careful consideration must be offered when designing solutions with this level of extensibility.
Because of the variability of the options available, there is no general architectural recommendation for handling configuration use cases in the ER system, nor are there generic abstractions which would be generally applicable to all ER systems, to the best of our knowledge. Following well established design principles and best practices is the strongest recommendation possible.
@@ 4.4 Entity Resolution Evaluation
Systems architectures should incorporate ER evaluation use cases. In terms of the recommended approach, there is no size which fits all available options.
@@ 5 ER Logical Components
The pipeline design pattern is a well-established concept in both literature and practice [67, 36]. It structures computation as a sequence of processing components, where each component receives input, transforms it, and passes the output to the next stage in the pipeline. This design pattern promotes modularity, configurability, and scalability, allowing each component to operate independently and process data as soon as it becomes available [67]. The pipeline structure closely mirrors neural network architectures, where layers sequentially process and refine representations. Prior work has successfully leveraged this structure for neural network-based ER systems [44], since ER involves sequential transformations of data, much like a pipeline.
@@ 5.1 Information Flow
The pipeline nature of entity resolution (ER) becomes evident when analysing its information flow at an abstract level. Figure 4 illustrates the runtime information flow for batch ER [12].
@@
This lack of an upper bound also applies to incremental ER, which differs from batch ER in two key aspects: the requirement for maintaining a runtime state and the continuous ingestion of input data rather than batch processing. Continuous ingestion also introduces the possibility of supporting new types of input sources, such as API endpoints.
@@
Regardless of the implementation details, entity resolution fundamentally operates as a pipeline, transforming input information records into entity profiles. The following subsections provide a detailed description of each component in Figures 4 and 5.
@@ 5.2 Entity Reference Extraction
Broadly, entity reference extraction follows two dominant paradigms. Some systems favor vector-based extraction[44, 33], producing numerical representations suitable for training or integrating with neural networks. More traditionally, attribute-based extraction[51, 34, 21, 28, 58] focuses on parsing structured semantic attributes from input data. A particularly relevant case of attribute-based extraction arises in Named Entity Recognition and Disambiguation (NERD) systems. Frameworks such as spaCy [42] and NLTK [7] extract named entities from unstructured text while also providing associated attributes. This structured output makes NERD-derived data structuring similar to that used in classical record linkage tasks.
A distinct case arises in entity alignment, where the extracted data is structured as a knowledge graph [70]. Unlike graph-based record linkage[58], which primarily stores similarity relationships between entity references, entity alignment operates over knowledge graphs, identifying correspondences across different graphs. While both methods generate graph representations, their treatment in the ER pipeline differs: entity alignment constructs a separate knowledge graph per data source, whereas graph-based record linkage represents each entity reference as a vertex[58], integrating all references into a single logical graph. This divergence from record linkage may be addressed by entity alignment system architectures at the level of the entity reference extraction component.
@@
Entity reference extraction may rely on abstractions named extraction traits. These are flexible, extensible algorithms that define the format, function, and meaning of extracted data. Traits can be simple, extracting fundamental data types like floating-point numbers or keywords, or compound, combining multiple traits. The main purpose of extraction traits is to adapt data representations to internal ones. Making extraction traits part of an ER system’s architecture allows tailoring its data ingestion to various use cases while maintaining a consistent pipeline for downstream ER tasks.
Despite methodological differences, all approaches ultimately produce an instance of an internal data structure specific to each ER system. The instance produced by an assortment of extraction traits is referred to as an entity reference.
@@ 5.3 Comparison Space Generator
Another observation on existing systems is that a reliance on both matching and clustering is not necessary for performing ER. Alternatives exist which bypass either matching[38] or clustering[44]. One might assume that omitting the traditional matching step removes the need to constrain the number of comparisons. However, the fundamental challenge of a quadratic comparison space remains in all cases.
@@
Figure 7 should be regarded as a straw man which highlights key mechanisms for controlling the comparison space within the ER runtime. Methods such as blocking, indexing, and filtering serve to reduce the number of comparisons that must be performed further in the ER pipeline. Herein, they are collectively referred to as comparison space generation methods.
A common architectural contract for designing comparison space generator observed across all reference systems assumes a sequence of entity references as input and entity references arranged in a way that minimizes the number of subsequent comparisons as output. Commonly, pairs of entity references are output because the common implementation of a comparison has two arguments. However, Figure 7 captures the posibility that multiple entity references participate in subsequent comparisons because of models [57] which generalize ER matching to this effect.
System architectures may wish to allow storing entity references from the comparison space in a dedicated reference store.
@@ 5.4 Matching Engine
While matching and clustering often work together in ER, they are not strictly dependent on each other. At least one must be present to resolve entities. Matching reflects the fundamental way in which an observer—human or algorithmic—identifies similarity between data points and prior knowledge. The presence of a matching subsystem in the ER pipeline has marked benefits. For example, a matching subsystem can produce a similarity graph which captures similarity relationships between entities [51]. Due to a lower amount of information presented to the user, interfaces highlighting individual pairs of potentially matching entity refernces are easier to build and more user friendly.
@@ 5.5 Reference Store
This component plays the central role in entity alignment where it is synonymous with either the source or the target knowledge graph [70, 63]. Even in simpler implementations, linked references represented as pairs of tuples are tracked using this component in-memory [21, 44, 33] or in Apache Spark tables [34].
Designs of this logical component are reminiscent of the Memento design pattern [19] because it is primarily used to store and access previous versions of the same data. Common concerns to consider include data availability and consistency.

### v2

@@ Abstract
Entity resolution (ER) is a critical task in data management which identifies whether multiple records refer to the same real-world entity. Despite its significance across domains such as healthcare, finance, and machine learning, implementing effective ER systems remains challenging due to the abundance of methodologies and tools, leading to a paradox of choice for practitioners. This paper proposes Resolvi, a reference architecture which prioritizes extensibility, interoperability, and scalability in ER systems. By analyzing existing ER frameworks and literature, we establish a structured approach to designing ER solutions that address common challenges. Additionally, we explore best practices for system implementation and deployment strategies to facilitate large-scale entity resolution. Through this work, we aim to provide a foundational blueprint that assists researchers and practitioners in developing robust, scalable ER systems while reducing the complexity of architectural decisions.
@@ 1 Introduction
The terminology surrounding ER is diverse. It is commonly referred to as record linkage, deduplication, merge/purge, or entity matching [12, 51, 6]. In some contexts, it appears under the names entity alignment or named entity recognition and disambiguation (NERD) [56]. This terminological variation reflects the breadth and complexity of the field, creating challenges in unifying methodologies and system architectures.
@@ 3.1 Business Requirements
The first such requirement is determining whether multiple information refer to the same real world object. With the risk of stating the obvious, this entails that the system should discern what a real world object is. Human assistance in this regard is arguably required at this point in our technological development. It logically follows that completely unsupervised entity resolution solutions are more unreliable than supervised ones. Consequently, humans should at least be allowed in the loop at design time.
@@ 3.2 Architectural Characteristics
Just as business requirements dictate whether a system performs entity resolution, certain architectural characteristics are universally required for systems that execute end-to-end entity resolution.
ER is a highly polymorphic task with a huge number of applications. Among the top priorities when implementing any such system is its extensibility. Extensibility is exhibited to various degrees in practice today. Perhaps the most important reason for making extensibility the top priority of this reference architecture is the variability of implementation choices made by the existing ER systems. Whether discussing a customizable number of data sources [60, 53] or the ability to completely replace the core runtime logic [30, 46], ER systems today are nothing if not flexible. The technological diversity ranging from systems using SQL [36] to ones using advanced neural networks [35] is another reason for the reference architecture to prioritize extensibility. This prioritization necessarily entails a higher degree of generalization and abstraction.
ER is valuable standalone or as part of a more complex process. Therefore, the best ER systems will be highly interoperable either by supporting many data contracts or by using widely adopted technologies and data formats. Interoperability is an obvious concern to some of the inspected reference systems [53, 21, 60, 36]. Because the reference architecture values interoperability, the concepts and recommendations presented favour the adoption of frameworks or standards wherever possible.
Basic ER implementations that rely on similarities between entity references have a minimum time complexity of $\mathcal{O}(n^{2})$ [12] where $n$ is the total number of entity references available to the system. The high complexity coupled with the need to process large amounts of data makes performance and scalability indispensable architectural characteristics of any modern ER system. Performance is the ability of a system to use the available hardware resources efficiently [55] (i.e the system does not waste computations). Scalability is the ability of a system to use hardware resources effectively as the demand on the system increases [55] (e.g the system is able to access more resources as it receives more input). As a consequence of valuing performance and scalability, the reference architecture insists on aspects which promote these qualities, such as blocking, filtering, parallel execution or distributed computation.
@@ 4.2 Entity Resolution Runtime
Some ER systems store entity references and entity profiles for later use. Storing the state is particularly important for performing progressive ER [60, 53, 21] or to preserve stable identifiers across time for entity profiles [70].
@@ 4.3 System Configuration
In existing systems, the end user can customize the runtime itself by choosing a default runtime [36, 60] or even building one from existing components exposed by the system [25, 53]. Another runtime customization is the ability to train a configuration optimised for a specific dataset either via active learning [68] or genetic algorithms [21]. This configuration can then be attached to many ER runtime executions. This approach stands in contrast to building the optimal configuration at runtime through unsupervised learning [36]. In that case a runtime configuration of the data record attributes that are relevant in ER is sufficient.
System configuration is exhibited in existing ER systems through plugin management for some (DeepMatcher [46] can use Magellan [30] for entity reference extraction) or for all of the ER runtime components (both FAMER [60] and JedAI [53] allow extending almost all components used at runtime). Some systems are designed as applications to be used by end users. JedAI [67] and FAMER [60] are in this category. These systems typically implement a runtime paradigm ((rule-based or graph-based, respectively)) which cannot be changed through configuration. In contrast, systems may be designed as frameworks [25, 30], enabling the configuration of the runtime architecture itself. The DeepMatcher design [46] allows using neural networks throughout the entire runtime or only a part of it by allowing the ER application developers to replace the way embeddings are generated, for example. This reference architecture favours framework design because it is more extensible. ER system designs should allow swapping out some or all runtime components at design time without changing the underlying abstractions presented within this reference architecture.
@@
Specific ER system architectures should involve administrators in most configuration use cases. This recommendation assumes that administrators are better prepared on security and configuration topics than end users and that in practice they usually have broader access to the system than developers. Developers also play a role in the design of the ER runtime. They may train machine learning models or implement plugins that perform various ER tasks such as matching, clustering or data ingestion. These models and plugins are visible to end users through configuration. Developers may register and unregister the models and plugins they author. They also typically maintain plugin implementations and models.
Plugin management may be performed by all three roles: administrators, end users and developers. End users might focus on choosing the plugins they need to configure custom runtimes, whereas developers would focus on managing their own plugins. Finally, administrators could enforce system-wide rules for using plugins. A system architecture that allows external developer involvement (through plugin management or otherwise), may be desirable. However, it raises many potential concerns, primarily from the sphere of data privacy and security. Careful consideration must be offered when designing solutions with this level of extensibility.
We saw that configuration can occur at three levels: parametrization, runtime configuration, and system configuration. The configuration of an ER system depends on its technical goal, problem domain, and technology context. This limits the reference architecture, as no universal recommendations or abstractions apply solely to ER systems. The best approach is to follow established design principles and best practices when defining logical components and the configuration processes. Throughout the remainder of this paper we will not insist on the configuration of ER systems.
@@ 4.4 Entity Resolution Evaluation
System architectures derived from this reference should incorporate ER evaluation use cases. In terms of a recommended logical or procedural approach, there is none. The reason behind this is that ER systems may be evaluated on design time or at runtime. Depending on the above choice, evaluation can be very simple (similar to what all existing systems implement today) or very complex (real-time monitoring of the results, for example). While ER evaluation design is important, it is adjacent to ER runtime design. Furthermore, reference implementations of ER evaluation systems which focus on the real-time evaluation of ER at runtime are missing, to the best of our knowledge. Throughout the remainder of this paper we will not dive deeper into topics related to ER evaluation, leaving that expansion on this paper as future work.
@@ 5 ER Logical Components
The pipeline design pattern is a well-established concept in both literature and practice [69, 38]. It structures computation as a sequence of processing components, where each component receives input, transforms it, and passes the output to the next stage in the pipeline. This design pattern promotes modularity, configurability, and scalability, allowing each component to operate independently and process data as soon as it becomes available [69]. The pipeline structure closely mirrors neural network architectures, where layers sequentially process and refine representations. Prior work has successfully leveraged this structure for neural network-based ER systems [46], since ER similarly involves sequential transformations of data.
@@ 5.1 Information Flow
The pipeline nature of entity resolution (ER) becomes evident when analysing its information flow at an abstract level. Figure 4 illustrates the runtime information flow for that which is termed batch ER in literature [12].
@@
This lack of an upper bound can lead to performing what is known as incremental ER [24, 12]. It differs from batch ER in two key aspects: the requirement for maintaining a runtime state and the continuous ingestion of input data rather than batch processing. Continuous ingestion also introduces the possibility of supporting certain architecture styles better (e.g service oriented architecture, event driven architecture).
@@
System Incremental Batch Source Code
ditto [35] + Magellan [30] ditto.run_all_er_magellan
deepmatcher [46] deepmatcher.runner
Duke duke.Duke
duke.server
JedAI Toolkit [53] jedai.workflowbuilder
FAMER [60] FamerConfiguration
IncrementalConfiguration
Splink [36] internals.database_api
internals.linker
d-blink [40] dblink.Run
Table 3: Incremental and batch capability of the reference ER systems.
Regardless of specific implementation details, entity resolution fundamentally follows a pipeline model, transforming input records into structured entity profiles. According to the original pipeline design pattern, the most effective approach to pipeline construction is component-based design. Figure 6 illustrates key top-level abstractions common to all reference ER systems, though they may appear in different forms in practice. The diagram emphasizes the diversity of input sources and the various ways output data can be represented. Additionally, it serves as a reminder that pipeline components primarily process entity references and entity profiles. This characteristic structure of the ER pipeline enhances system comprehensibility, particularly for newcomers. The key takeaway is that once entity references have been determined within each data source, they can be processed iteratively to generate entity profiles in the desired format using a series of pipeline components.
Figure 6: Top level ER abstractions which set the framework for ER, as described in Subsection 5.1.
The following subsections provide more detailed descriptions for key pipeline components from Figures 4 and 5.
@@ 5.2 Entity Reference Extraction
Broadly, entity reference extraction follows two dominant paradigms. Some systems favor extracting vectors from the input information, producing numerical representations suitable for training or integrating with neural networks. More traditionally, the extracted information is structured into objects with attributes. Note that such objects include node vertices, mapping types or columnar storage data types (e.g dataframes). Table 4 shows this entity reference extraction categorization for our selection reference systems.
System Representation Source Code
ditto [35] + Magellan [30] Attributes py_entitymatching.io
ditto_light.dataset
ditto_light.augmenter
deepmatcher [46] Attributes, py_entitymatching.io
Vectors, attr_summarizers
Hybrid
Duke Attributes datasources
N-Triples LuceneDatabase
JedAI Toolkit [53] Attributes jedai.datareader
jedai.datamodel
FAMER [60] Graph Vertices famer.preprocessing.io
Splink [36] Attributes splink.internals.database_api
splink.internals.splink_dataframe
d-blink [40] Attributes dblink.Project
Table 4: Entity reference extraction (described in Subsection 5.2) implementations in the reference systems.
Outside of our selection of reference ER systems, a particularly interesting case of attribute-based extraction arises in Named Entity Recognition and Disambiguation (NERD) systems. Frameworks such as spaCy [44] and NLTK [7] extract named entities from unstructured text while also providing associated attributes. This structured output makes NERD-derived data structuring similar to that used in classical record linkage tasks.
In the case of entity alignment the extracted data is structured as a knowledge graph [72]. Unlike graph-based record linkage implemented by FAMER[60], which primarily stores similarity relationships between entity references, entity alignment operates over knowledge graphs, identifying correspondences across different graphs. While both methods generate graph representations, their treatment in the ER pipeline differs: entity alignment constructs separate knowledge graphs per each data source, whereas graph-based record linkage represents each entity reference as a vertex, integrating all references into a single logical graph. This divergence from the typical record linkage logic must first be addressed by entity alignment system architectures in the entity reference extraction component.
@@
Entity reference extraction may rely on abstractions we refer to as extraction traits. These are flexible, extensible algorithms that define the format, function, and meaning of extracted data. Traits can be simple, extracting fundamental data types like floating-point numbers or keywords (e.g dblink.Project CSV parsing), or compound, combining multiple traits (e.g ditto_light.augment module which allows performing several chained operations on the data). The main purpose of extraction traits is to adapt data representations to internal ones. Making extraction traits part of an ER system’s architecture allows tailoring its data ingestion to various use cases while maintaining a consistent pipeline for downstream ER tasks.
Despite methodological differences, all approaches ultimately produce an instance of an internal data structure specific to each ER system. The instance produced by an assortment of extraction traits is referred to as an entity reference. The data structure useful in ER contains one or more entity references. An admitted limitation of the reference architecture is the inability to specify which data structure is most suitable. This limitation stems from the multiple available options, the choice between which is driven by the ER model choice. For example, in record linkage scenarios we might use an algebraic set whereas entity alingment requires using knowledge graphs.
@@ 5.3 Comparison Space Generator
Another observation on existing systems is that a reliance on both matching and clustering is not necessary for performing ER. Alternatives exist which bypass either matching[40] or clustering[46]. One might assume that omitting the traditional matching step removes the need to constrain the number of comparisons. However, the fundamental challenge of a quadratic comparison space remains in all cases, as pointed out by previous literature reviews [52].
@@
Figure 8 should be regarded as a straw man which highlights key mechanisms for controlling the comparison space within the ER runtime. Methods such as blocking, indexing, and filtering serve to reduce the number of comparisons that must be performed further in the ER pipeline.
System B I F Source Code
ditto [35] + Magellan [30] py_entitymatching.blocker
deepmatcher [46] py_entitymatching.blocker
Duke AbstractBlockingDatabase
LuceneDatabase
JedAI Toolkit [53] jedai.blockbuilding
jedai.blockprocessing
FAMER [60] famer.linking.blocking
linking.selection
Splink [36] splink.internals.blocking
d-blink [40] dblink.partitioning
Table 5: Comparison space generation techniques supported by the reference systems. B stands for blocking, I stands for indexing and F stands for comparison filtering.
Blocking, indexing and filtering are collectively referred to as comparison space generation methods from here on. Table 5 lists how each of the reference systems handles comparison space generation.
As with entity reference extraction, there is a limit to the amount of details that can be specified in the reference architecture. For example, while tempting to imagine modelling comparison space generation methods using a coherent unitary contract, this is not necessarily possible. The JedAI system provides a good example in this regard. Its org.scify.jedai.blockprocessing package defines several filtering and indexing implementations which expect as input the output of a blocking operation. Therefore, detailing the design of individual comparison space generators does not seem wise.
On the other hand, it is common in practice for pairs of entity references to be output by the comparison generation subsystem. This can be seen in practice across all reference systems, including FAMER (which outputs pairs of graph vertices). However, in Figure 8 we envision the posibility that multiple entity references participate in subsequent comparisons. The reference architecture is suitable for models which generalize ER matching to this effect [59] because of this. This provides an extension point for using such models.
Finally, implementing system architectures may wish to allow storing the comparison space in a dedicated reference store to reduce the running time of subsequent resolutions over the same data or to support incremental ER.
@@ 5.4 Matching Engine
While matching and clustering often work together in ER, they are not strictly dependent on each other. At least one must be present to resolve entities. Matching reflects the fundamental way in which an observer—human or algorithmic—identifies similarity between data points and prior knowledge. The presence of a matching subsystem in the ER pipeline has pragmatic benefits. For example, a matching subsystem can produce a similarity graph which captures similarity relationships between entities, resulting in more nuanced clustering strategies downstream [53, 60]. Additionally, due to a lower amount of information presented to the user, interfaces highlighting individual pairs of potentially matching entity refernces might be easier to build and easier to make user friendly.
@@ 5.5 Reference Store
This component plays the central role in entity alignment where it is synonymous with either the source or the target knowledge graph [72, 65]. For record linkage implementations, Duke’s Database components provide a flexible example of a reference store. Splink’s use of Apache Spark tables to store and quickly retrieve relevant data exemplified in its blocking module is another example which showcases how the reference store logical abstraction might be designed in practice.
When it is designed as a component, it is reminiscent of the Memento design pattern [19] because it is primarily used to store and access previous versions of the same data. Common concerns to consider when designing the reference store include data availability and integrity.